OpenStudy (mixhoneyy):

(1,1);y=x+12 Write an equation of line that passes through the given point and is parallel to the given line

TheSmartOne (thesmartone):

parallel lines have the same slope so find the slope from y = x + 12 y = mx + b, m is the slope in put the slope in to y = mx + b and then plug in (1, 1) in place of x and y respectively and solve for b and then write y = mx + b and plug in the slope and the value of b b is the y-intercept in y = mx + b show your work and I can check over it or guide you further :)